Show that the period of oscillation of main on helical spring is given by 

T = 2π √(k/m)



1
Expert's answer
2021-03-26T11:20:20-0400

The period of oscillations of a mass oh a helical spring depends on the spring constant and mass. According to Newton's second law:


"F=ma=m\frac{dx^2}{d^2t}."

On the other hand, this is equal to


"F=-kx."

So, we have


"m\frac{dx^2}{d^2t}=-kx."


We know that displacement depends on time according to the following equation:


"x=A\cos\bigg(\frac{2\pi t}{T}\bigg)."

Substitution in the equation according to Newton's second law gives


"-m\frac{(2\pi)^2}{T^2}A\cos\bigg(\frac{2\pi t}{T}\bigg)=-kA\cos\bigg(\frac{2\pi t}{T}\bigg),\\\space\\
m\frac{(2\pi)^2}{T^2}=k,\\\space\\
T=2\pi\sqrt{m/k}."